OpenStudy (anonymous):

Simplify (16x^12)^3/4

OpenStudy (campbell_st):

I'd rewrite 16 as 2^4 and the index law you need is power of a power \[(x^a)^b = x^{a \times b}\] so in your question your have \[(2^4x^{12})^{\frac{3}{4}} = (2^4)^{\frac{3}{4}} \times (x^{12})^{\frac{3}{4}}\] multiply the powers for the solution.. you may need to wwrite your power of 2 as a normal number

OpenStudy (anonymous):

So 8x^9

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

Thanx:-)